Exercise 1 Find and classify the stationary points of f(x, y) = (x² + y)e8/2.Exercise 2 A box with a volume of 8m3 is to be constructed with a gold-plated top, silver-plated bottom, and copper-plated si
Answer #1

1) The given function 9/ és $3,3) =(22+y).e =o 9/2 =0 9/2 2 = 0 2.e +y+2 2 x=0 let 25 = 0 & ab doc dy Y2 = 2xe I €(x2+4+2) =se, cat (0, 2) the function has relative minimce and minimum value es $40,-2) = -2 е 2) Let X , yo the dimension of the baseI You x = 1 x x3 = 1 x = 1 ㅠ ㅠ 7 2 y = 1 & Z = 8 box is: = 1 So, the dimension q wid length & height = 8 Hope this is helpful